Calculate the required probabilities for the normal distributions with the parameters specified in parts a through e.
a. μ = 5, σ = 4; calculate P(0 < x < 6).

P(0 < x < 6) = ______ . (Round to four decimal places as needed.)

b. μ = 5, σ = 5; calculate P(0 < x < 6).

P(0 < x < 6) = ______ . (Round to four decimal places as needed.)

c. μ = 4, σ = 4; calculate P(0 < x < 6).

P(0 < x < 6) = ______ . (Round to four decimal places as needed.)

d. μ = 6, σ = 2; calculate P(x > 1).

P(x > 1) = ______ . (Round to four decimal places as needed.)

e. μ = 0, σ = 2; calculate P(x > 1).

P(x > 1) = ______ . (Round to four decimal places as needed.)
Answer #1

Solution :

a.

P(0 < x < 6) = P[(0 - 5)/ 4) < (x - \mu ) /\sigma  < (6 - 5) / 4) ]

= P(-1.25 < z < 0.25)

= P(z < 0.25) - P(z < -1.25)

= 0.5987 - 0.1056

= 0.4931

P(0 < x < 6) = 0.4931

b.

P(0 < x < 6) = P[(0 - 5)/ 5) < (x - \mu ) /\sigma  < (6 - 5) / 5) ]

= P(-1 < z < 0.2)

= P(z < 0.2) - P(z < -1)

= 0.5793 - 0.1587

= 0.4206

P(0 < x < 6) = 0.4206

c.

P(0 < x < 6) = P[(0 - 4)/ 4) < (x - \mu ) /\sigma  < (6 - 4) / 4) ]

= P(-1 < z < 0.5)

= P(z < 0.5) - P(z < -1)

= 0.6915 - 0.1587

= 0.5328

P(0 < x < 6) = 0.5328

d.

P(x > 1) = 1 - P(x < 1)

= 1 - P[(x - \mu ) / \sigma < (1 - 6) / 2)

= 1 - P(z < -2.5)

= 1 - 0.0062

= 0.9938

P(x > 1) = 0.9938

e.

P(x > 1) = 1 - P(x < 1)

= 1 - P[(x - \mu ) / \sigma < (1 - 0) / 2)

= 1 - P(z < 0.5)

= 1 - 0.6915

= 0.3085

P(x > 1) = 0.3085